OpenStudy (anonymous):

PLEASE HELP ME WITH THIS PROBLEM If f(x)=lnx,g(x)=e^5x, and h(x)=x^8, find the following. (a) (f(g(x)) and what is the domain (b) g(f(x)) and what is the domain (c)(f(g(7)) (d)(f(h(x)) and what is the domain (e) (f(h(e))

jimthompson5910 (jim_thompson5910):

do you mean f(x) = lnx

OpenStudy (anonymous):

yes! thank you for catching that!

jimthompson5910 (jim_thompson5910):

f(x) = ln(x) f(g(x)) = ln(g(x)) f(g(x)) = ln(e^(5x)) f(g(x)) = 5x*ln(e) f(g(x)) = 5x

jimthompson5910 (jim_thompson5910):

the domain is the set of all real numbers since you can plug in any real number you want to get some real number as an output

OpenStudy (anonymous):

okay would it be the same for b? just change the number?

jimthompson5910 (jim_thompson5910):

well you'll get the same result, just take a different path

jimthompson5910 (jim_thompson5910):

g(x) = e^(5x) g(f(x)) = e^(5*f(x)) g(f(x)) = e^(5*ln(x)) g(f(x)) = e^(ln(x^5)) g(f(x)) = x^5 actually you won't get the same result, i jumped the gun a bit there...but the domain is still the set of all real numbers

OpenStudy (anonymous):

okay understand now.

jimthompson5910 (jim_thompson5910):

you're basically doing this for every part a) through e)

OpenStudy (anonymous):

would the next be equal to 35

jimthompson5910 (jim_thompson5910):

it would

jimthompson5910 (jim_thompson5910):

what about part d

OpenStudy (anonymous):

I got to the point where it is 8lnx not sure where to go from there

jimthompson5910 (jim_thompson5910):

that's as far as you can go

jimthompson5910 (jim_thompson5910):

so f(h(x)) = 8*ln(x) where the domain is x > 0

OpenStudy (anonymous):

oh! lol okay so then e will be just 8

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

is there a chance you can help me with just two more problems?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

that sounds doable

OpenStudy (anonymous):

Thank you SOoo much should I post it here or a new one?

jimthompson5910 (jim_thompson5910):

either way works